OpenStudy (anonymous):

how to write x^2+y^2=4x in terms of x=something

OpenStudy (tkhunny):

There are TWO answers. It will take either Completing the Square or the Quadratic Formula.

OpenStudy (anonymous):

You need to first put all the x-terms on one side and all the rest on the other side: x^2 + y^2 = 4x x^2 - 4x = -y^2 x(x-4) = -y^2 x = -y^2/(x-4)

OpenStudy (anonymous):

@tkhunny how do ya do it with quadratic formula

OpenStudy (anonymous):

@Rav that's not the way to roll, sorry :(

OpenStudy (tkhunny):

Just write it down. x^2+y^2=4x x^2 - 4x + y^2 = 0 \(x = \dfrac{4\pm\sqrt{(-4)^{2} - 4(1)(y^{2})}}{2}\) - Done

OpenStudy (anonymous):

OP, thx mate